# A Systems Thinking Approach to Algorithmic Fairness

### v3

@@ Abstract.
Systems thinking provides us with a way to model the algorithmic fairness problem by allowing us to encode prior knowledge and assumptions about where we believe bias might exist in the data generating process. We can then encode these beliefs as a series of causal graphs, enabling us to link AI/ML systems to politics and the law. This allows us to combine techniques from machine learning, causal inference, and system dynamics in order to capture different emergent aspects of the fairness problem. We can use systems thinking to help policymakers on both sides of the political aisle to understand the complex trade-offs that exist from different types of fairness policies, providing a sociotechnical foundation for designing AI policy that is aligned to their political agendas.
@@ 3. Towards a unified theory of fairness
We will use systems thinking to encode these different ideological beliefs into a series of causal graphs. This will help us to understand what types of interventions policymakers on both sides of the political aisle might want to make to ensure that a machine learning model is fair. Finally, this could give us insights into how to build machine learning models that are compliant with antidiscrimination laws.
@@ 4. Modeling fairness as a machine learning problem
Before we build a causal representation of the fairness problem, we will start by modeling fairness as a supervised machine learning problem using a simple block diagram as shown in Figure 1. In the center is a black box machine learning model that receives data $X$ as input and returns decision $D$ as output. The decision $D$ causes an outcome $Y$ that could potentially be recorded as future data $X$, thus creating a feedback loop. The key question that we need to address is whether the protected attribute $A$ should be used in a machine learning model and if so, how?
[…]
Figure 1. Block diagram representation of algorithmic fairness
[…]
A block diagram of a black box ML model receiving data $X$ as input and returning decision $D$ as output. Decision $D$ causes an outcome $Y$ which can become future data $X$. A protected attribute $A$ is being fed into the model.
@@ 5. Modeling fairness as a causal inference problem
We can transform the block diagram from Figure 1 into a causal Bayesian network (CBN) as shown in Figure 2. CBNs were pioneered by Pearl(Pearl, 2000; Pearl and Mackenzie, 2018; Pearl et al., 2016) and allow us to symbolically encode knowledge and assumptions about causal effects between variables, which are shown in this graph as nodes. The first node of the CBN represents a protected attribute $A$, which has a direct effect on a mediator $W$ ($A\rightarrow W$). The mediator $W$ has a direct effect on an outcome $Y$ ($W\rightarrow Y$). We assume that the mediator $W$ completely explains away the causal effect between the protected attribute $A$ and the outcome $Y$, resulting in conditional independence ($Y\perp\!\!\!\perp A|W$). However, the mediator $W$ might be considered biased by some people if they attribute the disparities across protected groups as being external. The mediator $W$ is filled white because it is latent, but it can still be partially measured as data $X$ ($W\rightarrow X$). Data $X$ is then fed into a machine learning model that is embedded inside the node for decision $D$ ($X\rightarrow D$). Finally, the decision $D$ also has a direct effect on an outcome $Y$ ($D\rightarrow Y$). CBNs are represented by directed acyclic graphs (DAGs), which means that we cannot use them to model feedback loops. Examples of what each of these variables could represent are shown in Table 1. For each of these applications, the protected attribute $A$ could represent race or gender for example.
[…]
Figure 2. Causal Bayesian network representation of algorithmic fairness
[…]
A causal Bayesian network where the protected attribute $A$ causes mediator $W$, which in turn causes an outcome $Y$ and is recorded as data $X$. The data $X$ is used to make a decision $D$ which can also affect outcome $Y$.
@@ 5.3. The limitations of causal Bayesian networks
CBNs have a key limitation: they can only be represented using directed acyclic graphs (DAGs). Recall from the block diagram in Figure 1 that outcome $Y$ could be recorded as future data $X$. We can modify the CBN from Figure 2 to model this feedback loop by adding a edge from outcome $Y$ to the mediator $W$ as shown in Figure 4 (this edge has been colored red), which is then recorded as data $X$. Even though this violates the DAG constraint of CBNs, we still need to model how different fairness policies may perpetuate disparities across protected groups over time. In the next section, we will show how to overcome this limitation using causal loop diagrams.
@@ 6. Modeling fairness as a system dynamics problem
We can transform the block diagram from Figure 1 and the CBN from Figure 2 into a causal loop diagram (CLD) as shown in Figure 5. CLDs were pioneered by Forrester (Forrester, 1971) and are similar to CBNs in that they allow us to symbolically encode knowledge and assumptions about cause and effect relationships between variables(Sterman, 2000). But unlike CBNs, CLDs allow us to model feedback loops. This is useful for modeling why we might believe that disparities continue to exist between protected groups.
[…]
Figure 5. Simplified causal loop diagram representation of algorithmic fairness
[…]
A causal loop diagram showing a decision $D$ being caused by data $X$ and the protected attribute $A$.
The CLD in Figure 5 is a simplified version of the full CLD that we will present later in this section. This shows how the same variables in the block diagram and the CBN can be mapped onto a CLD. In the center of the CLD is a node for decision $D$, which like before contains our machine learning model. This CLD is two sided, with the left side representing group A and the right side representing group B. Data from group A $X_{A}$ and group B $X_{B}$ is fed into the ML model, which renders a decision $D$ ($X_{A}\rightarrow D\leftarrow X_{B}$). Note that there is a time delay (denoted by ——) between decision $D$ and outcomes $Y_{A}$ and $Y_{B}$. The outcomes $Y_{A}$ and $Y_{B}$ are recorded as future data $X_{A}$ and $X_{B}$ ($Y_{A}\rightarrow X_{A}$ and $Y_{B}\rightarrow X_{B}$), which in turn drives a future decision $D$ ($X_{A}\rightarrow D\leftarrow X_{B}$). The outcomes $Y_{A}$ and $Y_{B}$ also influence whether the protected attribute $A_{A}$ or $A_{B}$ might be used to influence a decision $D$. For example, if group B has historically received worse outcomes that group A, then the group might demand the protected attribute $A_{B}$ to be affirmatively used to modify decision $D$ (fairness through affirmative action). However, the use of the protected attribute $A_{A}$ or $A_{B}$ could also impact outcomes $Y_{A}$ or $Y_{B}$.

### v4

@@ Abstract.
Systems thinking provides us with a way to model the algorithmic fairness problem by allowing us to encode prior knowledge and assumptions about where we believe bias might exist in the data generating process. We can then encode these beliefs as a series of causal graphs, enabling us to link AI/ML systems to politics and the law. This allows us to combine techniques from machine learning, causal inference, and system dynamics in order to capture different emergent aspects of the fairness problem. We can use systems thinking to help policymakers on both sides of the political aisle to understand the complex trade-offs that exist from different types of fairness policies, providing a sociotechnical foundation for designing AI policy that is aligned to their political agendas and with society’s values.
@@
[…]
Figure 1. Three representations of algorithmic fairness
[…]
A block diagram, a causal Bayesian network, and a causal loop diagram.
@@ 3. Towards a unified theory of fairness
We will use systems thinking to encode these different ideological beliefs into a series of causal graphs as shown in Figure 1. This will help us to understand what types of interventions policymakers on both sides of the political aisle might want to make to ensure that a machine learning model is fair. Finally, this could give us insights into how to build machine learning models that are compliant with antidiscrimination laws.
@@ 4. Modeling fairness as a machine learning problem
Before we build a causal representation of the fairness problem, we will start by modeling fairness as a supervised machine learning problem using a simple block diagram as shown in Figure 1a. In the center is a black box machine learning model that receives data $X$ as input and returns decision $D$ as output. The decision $D$ causes an outcome $Y$ that could potentially be recorded as future data $X$, thus creating a feedback loop. The key question that we need to address is whether the protected attribute $A$ should be used in a machine learning model and if so, how?
@@ 5. Modeling fairness as a causal inference problem
We can transform the block diagram from Figure 1a into a causal Bayesian network (CBN) as shown in Figure 1b. CBNs were pioneered by Pearl(Pearl, 2000; Pearl and Mackenzie, 2018; Pearl et al., 2016) and allow us to symbolically encode knowledge and assumptions about causal effects between variables, which are shown in this graph as nodes. The first node of the CBN represents a protected attribute $A$, which has a direct effect on a mediator $W$ ($A\rightarrow W$). The mediator $W$ has a direct effect on an outcome $Y$ ($W\rightarrow Y$). We assume that the mediator $W$ completely explains away the causal effect between the protected attribute $A$ and the outcome $Y$, resulting in conditional independence ($Y\perp\!\!\!\perp A|W$). However, the mediator $W$ might be considered biased by some people if they attribute the disparities across protected groups as being external. The mediator $W$ is filled white because it is latent, but it can still be partially measured as data $X$ ($W\rightarrow X$). Data $X$ is then fed into a machine learning model that is embedded inside the node for decision $D$ ($X\rightarrow D$). Finally, the decision $D$ also has a direct effect on an outcome $Y$ ($D\rightarrow Y$). CBNs are represented by directed acyclic graphs (DAGs), which means that we cannot use them to model feedback loops. Examples of what each of these variables could represent are shown in Table 1. For each of these applications, the protected attribute $A$ could represent race or gender for example.
@@ 5.3. The limitations of causal Bayesian networks
CBNs have a key limitation: they can only be represented using directed acyclic graphs (DAGs). Recall from the block diagram in Figure 1a that outcome $Y$ could be recorded as future data $X$. We can modify the CBN from Figure 1b to model this feedback loop by adding a edge from outcome $Y$ to the mediator $W$ as shown in Figure 3 (this edge has been colored red), which is then recorded as data $X$. Even though this violates the DAG constraint of CBNs, we still need to model how different fairness policies may perpetuate disparities across protected groups over time. In the next section, we will show how to overcome this limitation using causal loop diagrams.
@@ 6. Modeling fairness as a system dynamics problem
We can transform the block diagram from Figure 1a and the CBN from Figure 1b into a causal loop diagram (CLD) as shown in Figure 1c. CLDs were pioneered by Forrester (Forrester, 1971) and are similar to CBNs in that they allow us to symbolically encode knowledge and assumptions about cause and effect relationships between variables(Sterman, 2000). But unlike CBNs, CLDs allow us to model feedback loops. This is useful for modeling why we might believe that disparities continue to exist between protected groups.
The CLD in Figure 1c is a simplified version of the full CLD that we will present later in this section. This shows how the same variables in the block diagram and the CBN can be mapped onto a CLD. In the center of the CLD is a node for decision $D$, which like before contains our machine learning model. This CLD is two sided, with the left side representing group A and the right side representing group B. Data from group A $X_{A}$ and group B $X_{B}$ is fed into the ML model, which renders a decision $D$ ($X_{A}\rightarrow D\leftarrow X_{B}$). Note that there is a time delay (denoted by ——) between decision $D$ and outcomes $Y_{A}$ and $Y_{B}$. The outcomes $Y_{A}$ and $Y_{B}$ are recorded as future data $X_{A}$ and $X_{B}$ ($Y_{A}\rightarrow X_{A}$ and $Y_{B}\rightarrow X_{B}$), which in turn drives a future decision $D$ ($X_{A}\rightarrow D\leftarrow X_{B}$). The outcomes $Y_{A}$ and $Y_{B}$ also influence whether the protected attribute $A_{A}$ or $A_{B}$ might be used to influence a decision $D$. For example, if group B has historically received worse outcomes that group A, then the group might demand the protected attribute $A_{B}$ to be affirmatively used to modify decision $D$ (fairness through affirmative action). However, the use of the protected attribute $A_{A}$ or $A_{B}$ could also impact outcomes $Y_{A}$ or $Y_{B}$.